### v2

# Linguistic Nepotism: Trading-off Quality for Language Preference in Multilingual RAG

Abstract
Multilingual Retrieval-Augmented Generation (mRAG) systems enable language models to answer knowledge-intensive queries with citation-supported responses across languages. While such systems have been proposed, an open questions is whether the mixture of different document languages impacts generation and citation in unintended ways. To investigate, we introduce a controlled methodology using model internals to measure language preference while holding other factors such as document relevance constant. Across eight languages and six open-weight models, we find that models preferentially cite English sources when queries are in English, with this bias amplified for lower-resource languages and for documents positioned mid-context. Crucially, we find that models sometimes trade-off document relevance for language preference, indicating that citation choices are not always driven by informativeness alone. Our findings shed light on how language models leverage multilingual context and influence citation behavior. [^4] Code and data are released at https://github.com/dayeonki/linguistic_nepotism.
1 Introduction
Retrieval-Augmented Generation (RAG) systems have become a core component of modern large language model (LLM) pipelines, enabling models to answer knowledge-intensive queries by supplementing their limited parametric knowledge with external information (Lewis et al., 2020; Karpukhin et al., 2020; Gao et al., 2024). Given that over 50% of digital content is produced in languages other than English (Statista, 2025), recent work has extended these systems to multilingual RAG (mRAG) settings, which handle queries and documents in languages beyond English (Chirkova et al., 2024; Wu et al., 2024).
Despite recent advances, prior work highlights a key challenge in mRAG systems: language preference – a systematic tendency of models to favor sources written in certain languages during generation (Park & Lee, 2025). Understanding this behavior is crucial, as citation patterns shape both the information users see and the languages prioritized in multilingual knowledge access.
Existing approaches to measuring language preference, however, often fail to capture citation correctness. In short-form mRAG, preference has been estimated via information overlap (Sharma et al., 2025) or embedding similarity (Park & Lee, 2025), which do not directly account for correctness. In long-form mRAG, where outputs contain in-line citations (Zheng et al., 2025; Xu & Peng, 2025), preference has typically been measured by comparing citation frequencies against the language distribution of retrieved documents. This signal is coarse and confounded by the relevance and informativeness of multilingual sources ($\mathrm{C_{1}}$). Moreover, in-line citations are prone to hallucinations (Gao et al., 2023; Zhang et al., 2024), making it unclear whether observed preferences reflect true attribution or spurious citations ($\mathrm{C_{2}}$).
To address these challenges, we propose a controlled methodology for measuring language preference using model internal metrics (illustrated in Figure 1). We first construct a synthetic multi-parallel dataset of relevant documents, which allows us to isolate the effect of language while controlling for other factors such as document content and relevance (Step 1+2; addresses $\mathrm{C_{1}}$). Citation correctness is then verified through a two-step filtering process (Step 3; addresses $\mathrm{C_{2}}$) (§3.1). Next, we compare the accuracy of next token citation predictions (e.g., predicting “2” for document ID 2) while varying the language of the same cited document and keeping other variables fixed, including the language of remaining documents, document positions in the input context, and the query language (Step 4). Differences in citation accuracy between languages indicate a preference for the higher-accuracy language (§3.2).
Figure 1: Overview of our approach for measuring language preference. We show both synthetic data generation and measurement method. Given an English query $q$ and its $K$ relevant evidence documents $D_{en}$, we first translate the documents into multiple languages $D_{fr},D_{ar},D_{ko}\dots$ (Step 1). We then generate a reference citation-supported report $r$ for each query using $q$ and $D_{en}$ (Step 2). The report $r$ consists of sentence-level statements $s_{i}$, each paired with a single citation ID $c_{i}$. For each $r$, we retain only statements that are verified (Step 3). Language preference is detected when the next token prediction accuracy for the correct citation ID decreases as the language of the cited document is varied (Step 4).
Using this setup across eight languages and six open-weight models, we address the overarching question: Do models preferentially cite documents in certain languages during long-form mRAG? To further inform building more robust mRAG systems, we empirically address three key questions: (a) What factors amplify language preference? (b) What role does the query language play in language preference? and (c) Is citation behavior driven more by document relevance or language?
Our main findings can be summarized as follows:
  • Evidence of strong English preference: Across all tested models, we find a pronounced tendency to cite English documents when the query is in English. This preference amplifies when: (1) the cited document is in a lower-resource language (e.g., Bengali, Swahili), or (2) the cited document appears in the middle of the input context (§5).
  • Language preference towards query language: We show that language preference extends beyond English: models favor citing evidence documents written in the query language (§6).
  • Language outweigh relevance: Last but not least, we show that models frequently cite English documents even when they are irrelevant to the query, suggesting that language itself exerts a stronger influence than document relevance in long-form mRAG (§7).
2 Related Work
Multilingual RAG. A growing body of work has examined that large language models (LLMs) are prone to hallucinations, especially in knowledge-intensive tasks (Augenstein et al., 2024; Huang et al., 2025a). Retrieval-augmented generation (RAG) mitigates this by retrieving external knowledge sources and incorporating them into generation (Chen et al., 2024; Gao et al., 2024). While early RAG systems largely focused on processing English queries and sources, recent research has extended these methods to multilingual RAG (mRAG), enabling retrieval and generation across a wider range of languages (Asai et al., 2022). Prior mRAG studies primarily examine the effects of query language (Chirkova et al., 2024), the language of relevant or irrelevant evidence documents (Wu et al., 2024; Qi et al., 2025; Liu et al., 2025), document ordering (Ranaldi et al., 2025a), and prompting strategies (Ranaldi et al., 2025b) on performance. However, due to cost efficiency and scalability (Saad-Falcon et al., 2024; Es et al., 2024), most of this work targets short-form mRAG, where the output is a brief answer to a factoid-style query (e.g., What is the capital of France?). In contrast, we focus on long-form mRAG, where models are asked to generate citation-supported reports in response to open-ended queries (e.g., How does fingerprint unlock work on phones?).
Long-form (m)RAG. Long-form RAG systems build upon prior work on long-form question answering (LFQA) datasets (Dasigi et al., 2021; Stelmakh et al., 2023) to generate paragraph level, citation-supported responses for complex, knowledge-intensive queries (Zhao et al., 2024; Wei et al., 2024; Ju et al., 2025; Zhang et al., 2025). Although evaluating models on long-form outputs is notoriously challenging (Qi et al., 2024), it is also increasingly important as it better mirrors how humans naturally interact with search engines (Khashabi et al., 2021), making such systems more easily integrable into search-based workflows like Deep Research platforms (Huang et al., 2025b; Zheng et al., 2025). Similarly, we use a long-form RAG dataset, Explain Like I’m Five (ELI5) (Fan et al., 2019), to measure language preference.
Language Preference. Language preference describes a systematic tendency for models to favor sources in certain languages over others. This preference largely arises from differences in training data distribution, tokenization methods, and resource availability (Wu et al., 2024; Sharma et al., 2025; Shen et al., 2024). Such preference manifests at both the retrieval and generation stages. On the retrieval side, prior work shows that multilingual information retrieval (MLIR) systems tend to favor high-resource languages (e.g., English) while under-representing sources in lower-resource languages, which can degrade retrieval quality (Telemala & Suleman, 2022; Yang et al., 2024; Amiraz et al., 2025) and introduce inconsistencies in generation (Chataigner et al., 2024). On the generation side, language models have been found to more effectively utilize sources written in specific languages (Park & Lee, 2025). Existing studies on short-form mRAG typically measure this by querying models in different languages and analyzing information overlap (Sharma et al., 2025) or embedding similarity (Park & Lee, 2025) between outputs and reference answers. In the long-form setting, prior work approximates language preference by comparing citation rates against the distribution of available documents per language, where over-representation in citations signals bias (Li et al., 2025). We build our work on this line of measuring language preference in long-form mRAG, but through a more controlled experimental setup using model internal metrics.
3 Measuring Language Preference in Long-form mRAG
Our goal is to measure whether LLMs systematically prefer citing evidence in some languages over others. To do this, we need (a) a multilingual dataset of queries with parallel evidence documents and verifiable citation-supported reports (§3.1), and (b) a measurement method that compares citation accuracy when the same document is presented in different languages (§3.2). Figure 1 shows the pipeline for dataset construction and measurement. All prompts are provided in Appendix A.
3.1 Synthetic Data Generation
Step 1: Evidence Document Translation. Let $\mathcal{D}_{en}=\{d_{1},\dots,d_{K}\}$ denote the set of $K$ relevant evidence documents in English associated with a query $q$. Since no parallel long-form mRAG datasets are publicly available, we construct multilingual variants $\mathcal{D}_{\ell_{\mathrm{target}}}$ for each target language $\ell_{\mathrm{target}}\in\mathcal{L}_{\mathrm{target}}$ using machine translation (MT). If $\mathrm{MT}_{\ell}$ denote a translation function into language $\ell$, we obtain $\mathcal{D}_{\ell}=\{\mathrm{MT}_{\ell}(d_{1}),\dots,\mathrm{MT}_{\ell}(d_{K})\}$. In our experiments, $\mathrm{MT}_{\ell}$ is implemented using Google Translate API. Despite the challenges of translating long-context documents (Wang et al., 2023; Cui et al., 2024; Wang et al., 2025b), the translation quality remains reasonable, with average COMET [^5] Unbabel/wmt22-cometkiwi-da quality estimation scores of 0.541. Per-language scores are reported in Appendix D.1.
Step 2: Reference Report Generation. For each query $q$ with associated English evidence document set $\mathcal{D}_{en}=\{d_{1},\dots,d_{K}\}$, we generate a reference citation-supported report using a strong LLM $\mathcal{M}_{\mathrm{gen}}$. We select OpenAI o3666https://openai.com/index/introducing-o3-and-o4-mini/ as $\mathcal{M}_{\mathrm{gen}}$, since its outputs were rated highest by human evaluators in SciArena (Zhao et al., 2025), a benchmark assessing long-form report generation and citation quality. The generated report is: $r=\mathcal{M}_{\mathrm{gen}}(q,\mathcal{D}_{en})$. [^7] On average, reports contain 148.5 words across 4.90 sentences. We segment $r$ into $n$ sentence-level statements: $r=(s_{1},[c_{1}],\dots,s_{n},[c_{n}])$, where $s_{i}$ is the $i$-th statement, and $c_{i}\in\{1,\dots,K\}$ is the citation ID of the evidence document $d_{c_{i}}\in\mathcal{D}_{en}$ that $\mathcal{M}_{\mathrm{gen}}$ cites as supporting $s_{i}$. By construction, $c_{i}$ denotes the citation token appearing in the report after $s_{i}$.
Step 3: Statement Pool Construction. Long-form generation with citations is prone to hallucination, with LLMs often introducing factual errors (Ji et al., 2023) or misattributing information to incorrect evidence (Gao et al., 2023; Magesh et al., 2024; Zhang et al., 2024). To ensure that only verifiably supported statements are retained for evaluation, we apply a two-stage filtering pipeline to the set of statement-citation pairs $\{(s_{i},c_{i})\}^{n}_{i=1}$ from Step 2. We perform filtering only if $|c_{i}|=1$ (i.e., statements with exactly one citation). First, the LLM-as-Relevance-Judge identifies statements whose cited document is deemed most relevant by the majority of judges. [^8] Prior work shows that LLMs provide precise relevance assessments (Ma et al., 2023; Sun et al., 2023). Second, the NLI entailment check verifies that the cited document actually entails the information in the statement.
(1) LLM-as-Relevance-Judge: Let $\mathcal{M}_{\mathrm{judge}}=\{m_{1},m_{2},m_{3}\}$ be the set of judge models that rank highest on the SciArena benchmark (OpenAI o4 mini6, Qwen-3 32B (Yang et al., 2025), and Gemini 2.5 Pro [^9] https://deepmind.google/models/gemini/pro/). Each judge $m\in\mathcal{M}_{\mathrm{judge}}$ is prompted with statement $s_{i}$ and the full evidence document set $\mathcal{D}_{en}$ to return the index of the most relevant document $j_{m}(s_{i},\mathcal{D}_{en})$. Here, $j_{m}$ implements a relative selection task over all $\mathcal{D}_{en}$ (i.e., “Which document best supports the statement?”), rather than an absolute binary support judgment (i.e., “Does this document support the statement?”), following findings that comparative framing improves LLM evaluation accuracy (Godfrey et al., 2025; Shrivastava et al., 2025). The total number of judges selecting the cited document $d_{c_{i}}$ is:
$\mathrm{\mathbf{votes}}(s_{i},c_{i})=\sum_{m\in\mathcal{M}_{\mathrm{judge}}}{\mathbbm{1}(j_{m}(s_{i},\mathcal{D}_{en})=c_{i})}$ (1)
We retain $s_{i}$ if when the majority of judges agree on the correct judgment: $\mathrm{\mathbf{votes}}(s_{i},c_{i})\geq 2$.
(2) NLI Entailment: We use an off-the-shelf Natural Language Inference (NLI) classifier $\phi$(premise, hypothesis) [^10] MoritzLaurer/mDeBERTa-v3-base-xnli-multilingual-nli-2mil7, which outputs 1 if the premise entails the hypothesis, and 0 otherwise. In our setting, $d_{c_{i}}$ is the premise and $s_{i}$ the hypothesis. We retain $s_{i}$ if $\phi(d_{c_{i}},s_{i})=1$. This is in accordance with the Attributable to Identified Sources (AIS) framework (Rashkin et al., 2023).
In practice, the LLM-as-Relevance-Judge and NLI Entailment filtering stages achieve retain rates of 90.35% and 96.12%, respectively. The final pool consists of 792 statements that pass both filters [^11] On average, each verified statement contains 33.70 words., ensuring that the correctness of each citation used for evaluation is reliably verified. [^12] Human annotation results in Appendix C show high agreement with the automatic filtering judgments.
3.2 Measurement Method
Step 4: Next Token Prediction Analysis. Intuitively, if the model predicts the correct citation token when the cited document is in English compared to another language, this indicates a preference for English. To quantify this, for each verified statement-citation pair $(s_{i},c_{i})$, we measure the accuracy of the model predicting $c_{i}$ as the top-1 next token.
We first construct a citation prediction prompt ending in the form: $x_{i}=s_{i}$[, where the bracket [ signals the start of the citation. To test for language preference for English, we define the set of evaluation languages as $\mathcal{L}_{\mathrm{eval}}=\{\mathrm{en}\}\cup\mathcal{L}_{\mathrm{target}}$, which includes English and all target languages. For each statement, we construct contrastive contexts where only the document to be cited, $d_{c_{i}}$, is presented in a language $\ell\in\mathcal{L}_{\mathrm{eval}}$, while all other evidence documents remain in English. The full context is denoted as $\mathrm{Context}(d_{c_{i}}\rightarrow\ell,d_{\neg c_{i}}\rightarrow\mathrm{en})$. Given the prompt prefix $x_{i}$, the model’s next token probability of the correct citation ID token $c_{i}$ corresponding to document $d_{c_{i}}$ conditioned on this context is: $p_{\theta}^{(\ell)}(c_{i})=\mathcal{P}_{\theta}(t=c_{i}|x_{i},q,\mathrm{Context}(d_{c_{i}}\rightarrow\ell,d_{\neg c_{i}}\rightarrow\mathrm{en}))$, where $\mathcal{P}$ is the model’s next token distribution given a prefix, and $\theta$ denotes model parameters. We define the model’s top-predicted citation token as: $\hat{c}^{(\ell)}_{i}=\mathrm{argmax}_{t}(p_{\theta}^{(\ell)})$, and compute citation accuracy in language $\ell$ over $n$ statements as:
$\mathbf{Acc}^{(\ell)}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(\hat{c}^{(\ell)}_{i}=c_{i}).$ (2)
A model exhibits English preference over a target language $\ell_{\mathrm{target}}\in\mathcal{L}_{\mathrm{target}}$ if it achieves higher citation accuracy when the cited document $d_{c_{i}}$ is in English than when it is in the target language. We define the citation accuracy gap as:
$\Delta(\ell_{\mathrm{target}})=\mathrm{\mathbf{Acc}}^{(\ell_{\mathrm{target}})}-\mathrm{\mathbf{Acc}}^{(\mathrm{en})}.$ (3)
In other words, $\Delta(\ell_{\mathrm{target}})$ quantifies how much more accurately the model cites English documents compared to the target language, with all other documents fixed to English. To ensure differences in raw scores are statistically meaningful, we perform pairwise two-sided $t$-tests and apply a Bonferroni correction to account for multiple comparisons.
4 Experiment Setup
Dataset. We use ELI5 dataset (Fan et al., 2019) of long-form questions from the Reddit forum “Explain Like I’m Five”. For each query, we adopt the WebGPT test set (Nakano et al., 2022) (270 queries), with relevant evidence documents collected by human annotators using Bing. To successfully answer a query, the generated output must cite all provided relevant documents. To ensure the citation IDs are tokenized as single tokens across all evaluated models, we only use queries with $K<10$ evidence documents. Detailed dataset statistics are in Appendix Table 2.
Languages. For $\mathcal{L}_{\mathrm{target}}$, we study eight languages representing a diverse range of resource levels (measured by number of speakers and Wikipedia articles), language families, scripts, linguistic typologies: Arabic (ar), Bengali (bn), Spanish (es), French (fr), Korean (ko), Russian (ru), Swahili (sw), and Chinese (zh). Detailed characteristics per language are outlined in Appendix Table 3.
Models. We use six open-weight LLMs that provide full-access to model weights and support large enough context windows to handle long-context evidence documents and long-form generations. To assess the generality of language preference, we evaluate models varying in size, degree of multilinguality, and architecture family: LLaMA-3.1 8B and LLaMA-3.3 70B (Grattafiori et al., 2024), Qwen-3 8B and 14B (Yang et al., 2025), Gemma-3 27B (Team et al., 2025), and Aya23 8B (Aryabumi et al., 2024). Details for each model can be found in Appendix Table 4.
5 Evidence of an English Language Preference
We seek to understand whether models prefer citing evidence documents in English over other languages in long-form mRAG. To do so, we analyze language preference in a controlled setup where all provided evidence documents are relevant to the query. We begin by comparing citation accuracies across languages, then explore factors that may impact language preference (§5.1). Next, we perform a layer-wise analysis of model behavior to unfold how language preference evolves (§5.2).
5.1 Do Models Preferentially Cite English Documents?
We define a model exhibits language preference for citing English evidence over the target language if its citation accuracy is higher for English ($\Delta(\ell_{\mathrm{target}})<0$ in Eq. 3). Table 1 presents citation accuracies by model and language. Overall, we see a consistent English preference across all tested models and target languages. Even models explicitly trained on diverse languages and multilingual tasks, such as Aya23 8B, display this preference. In Appendix D.2, we further show that, for all models, the next token probability of the correct citation ID is the highest and the Shannon entropy of the next token distribution is the lowest when the cited document is in English, indicating models are not only more accurate but also more confident in their correct predictions when to English evidence documents. We also find that smaller models (8B) have lower English baseline accuracy than larger models (e.g., LLaMA-3.1 70B, Gemma-3 27B), suggesting that models’ general ability to correctly cite English evidence documents tends to improve with model scale.
Language LLaMA-3.1 8B Qwen-3 8B Aya23 8B Qwen-3 14B Gemma-3 27B LLaMA-3.3 70B
English 67.4 62.6 60.0 83.0 86.2 85.9
French 62.9 (-4.49) 48.4 (-14.2)*** 48.5 (-11.5)*** 76.0 (-7.04)*** 79.0 (-7.21)** 77.4 (-8.50)***
Russian 62.1 (-5.30)* 50.4 (-12.2)*** 48.1 (-11.9)*** 74.8 (-8.17)*** 77.1 (-9.12)*** 74.5 (-11.4)***
Spanish 62.1 (-5.32)* 51.9 (-10.7)*** 49.1 (-10.9)*** 77.4 (-5.61)* 80.2 (-6.04)** 76.0 (-9.90)***
Korean 61.7 (-5.68)* 49.7 (-12.9)*** 42.2 (-17.8)*** 70.3 (-12.7)*** 77.5 (-8.71)*** 69.2 (-16.7)***
Chinese 59.9 (-7.51)* 49.2 (-13.4)*** 46.3 (-13.7)*** 73.5 (-9.49)*** 75.4 (-10.8)*** 74.1 (-11.8)***
Arabic 59.5 (-7.91)** 47.6 (-15.0)*** 43.2 (-16.8)*** 72.6 (-10.4)*** 78.4 (-7.82)*** 67.3 (-18.6)***
Bengali 56.6 (-10.8)*** 41.3 (-21.3)*** 27.2 (-32.8)*** 65.4 (-17.6)*** 77.9 (-8.33)*** 68.8 (-17.1)***
Swahili 53.0 (-14.4)*** 30.4 (-32.2)*** 22.4 (-37.6)*** 54.7 (-28.3)*** 74.0 (-12.2)*** 67.3 (-18.6)***
Table 1: Citation accuracies (%) by model and language. We present mean accuracy values $\mathrm{\textbf{Acc}}^{(\ell)}$ with $\Delta(\ell_{\mathrm{target}})$ in subscript. Pairwise two-sided $t$-tests are performed to compare accuracy between English and the target language, with the null hypothesis that the mean citation accuracy is equal across languages. Bonferroni correction is applied for multiple comparisons. *: significant with $p$ << 0.05; **: $p$ << 0.01; ***: $p$ << 0.001; non-marked: not statistically significant. Color coding indicates the magnitude of $\Delta(\ell_{\mathrm{target}})$: largest, second largest, others. Columns: increasing model size; rows: decreasing $\Delta(\ell_{\mathrm{target}})$ (of first model). All models consistently show English preference.
Stronger English Preference over Lower-resource Languages. Having established an overall preference for citing English documents, we next examine which factors amplify this preference. Using the $\Delta(\ell_{\mathrm{target}})$ values from Table 1 (i.e., the drop in citation accuracy relative to English), we find a clear correlation with language resource level: lower-resource languages exhibit largest accuracy decreases. For example, Swahili shows the greatest drop (-23.9% on average, up to -37.6% in Aya23 8B), followed by Bengali (-18.0% on average, up to -32.8% in Aya23 8B), even for models that officially support these languages (Qwen-3 8B, 14B, Gemma-3 27B; Appendix Table 4). In contrast, higher-resource languages such as Spanish and French show smaller decreases (-8.08% and -8.82% on average, respectively), indicating weaker English preference.
Figure 2: English accuracy (left) and the average of $\Delta(\ell_{\mathrm{target}})$ (right) (%) binned by relative position. Each bin is normalized by sample size. $\Delta(\ell_{\mathrm{target}})$ is largest when the cited document is positioned in the middle, indicating that position bias further amplifies English preference.
Position Bias Amplifies Language Preference. We find that the relative position of an evidence document within the input context impacts citation accuracy. Figure 2 (left) shows English citation accuracy binned by the relative position of the cited document: at the beginning (First), the end (Last), or elsewhere (Middle) in the input context. Accuracy is generally lowest when the document appears in the middle (one exception is LLaMA-3 70B, which shows the lowest accuracy for the Last position). This aligns with the “lost in the middle” phenomenon, where LLMs struggle to access and use information in the middle of long contexts (Liu et al., 2024), here demonstrated for citation generation. Figure 2 (right) presents the difference in accuracy between English and the average of target languages across these positions. [^13] Results for each target language can be found in Appendix D.3. For all models, the largest drop in accuracy occurs when the cited document is positioned in the middle, indicating that document position not only impacts English accuracy but also amplifies models’ English preference.
In sum, we provide strong evidence that models preferentially cite English evidence documents over target languages. This finding holds not only for corroborative attribution, which identifies sources that support a statement, but also for contributive attribution, which captures sources that causally influence the model’s generation, showing consistent trends (Appendix E). We further identify two key factors that amplify this preference: the resource level of the language and the position of the document within the input context.
5.2 Model Layer-wise Analysis
While our earlier results confirm a strong English preference in citation, we still lack a deeper understanding on how this preference unfolds during generation. Does the model settle on its initial choice and persist with it or does it initially favor English documents before shifting toward the correct target language citation? This question extends prior findings from short-form tasks, where multilingual LLMs often align their internal representations with English in early layers, transitioning to target language-specific spaces only in the final layers (Wendler et al., 2024; Zhong et al., 2024; Wang et al., 2025a; Bafna et al., 2025; Schut et al., 2025). We ask whether citation generation in long-form setup follows a similar trajectory: do models initially gravitate toward citing English documents and only later correct themselves, or is the outcome largely decided as soon as the model chooses which document to cite?
To probe this, we employ logit lens (Nostalgebraist, 2020), which maps intermediate state representations of LLMs into the vocabulary space, enabling the ability to track a model’s token prediction across layers. Since logit lens is tailored to probe a single token, our citation format is a single digit, and this approach works well for this use case. For each statement, we check whether the top-1 token prediction at a given layer is (1) the correct citation ID $c_{i}$ (target language document), (2) an incorrect ID $c_{j}$ ($j\neq i$, English document), or (3) not a valid citation token ($\notin\{1,\dots,K\}$, Others).
Figure 3: Logit lens visualization per language for LLaMA-3.1 8B (32 layers). $x$-axis: Last layer index; $y$-axis: Statement count. : Correct citation ID of document in target language; : Incorrect citation ID of document in English; : Not in valid citation set. Model makes a specific decision point when selecting which document to cite and largely preserves this choice across later layers. We only show last 14 layers. Results for other models are in Appendix D.4.
Figure 3 shows results for LLaMA-3.1 8B. Across all languages, layers 1-17 yield no valid predictions, indicating that the model has not yet figured out the expected output format. Around layers 18-20, both correct and incorrect citation IDs begin to appear, with correct IDs slightly more frequent. Layer 22 marks a sharp peak for both correct and incorrect predictions, suggesting this is the stage where the model settles on the output format and decides which document to cite. From layer 23 onward, incorrect IDs remain at a stable rate, showing that once the model commits to an incorrect citation, it rarely changes. Meanwhile, count for correct IDs steadily increases, replacing the earlier invalid predictions (Others). We also find that the gap between correct and incorrect predictions narrows notably for lower-resource languages (i.e., Bengali, Swahili), confirming our earlier findings that these languages exhibit a stronger English preference. Results for other models are provided in Appendix D.4.
Overall, these results indicate that models do not initially favor citing incorrect English documents and then switch to the correct target language. Instead, there is a specific decision point (around layer 22 for LLaMA-3.1 8B), when they decide which document to cite. From that point on, they largely preserve their initial decision, whether it is correct or not.
Figure 4: Accuracy per model for queries in the target language. : All docs in query language; : Cited document in English, all other docs in query language; : All docs in English; : Cited document in query language, all other docs in English. Note that $y$-axis scale vary by model. $x$-axis denotes different choices of target language. Models generally show query language preference.
6 Effect of the Query Language
Our previous analysis demonstrate that models preferentially cite English evidence documents over those in the target language. A natural follow-up question is whether this pattern persists when the query itself is in a language other than English: do models still prefer English documents, or do they prefer documents in the same language as the query?
Setting. We follow the same procedure used to measure English preference (Section 3), with one modification in Step 2 (reference report generation). Each user query is translated into the target language $q_{tgt}$, and for each, we generate a reference citation-supported report $r_{tgt}$ using $K$ relevant evidence document translations $\mathcal{D}_{tgt}$. [^14] We use Google Translate API for query translation, with translation quality reported in Appendix Table 5. For Step 4 (next token prediction analysis), we consider four context variants differing in the language of the cited document $d_{c_{i}}$ and the remaining evidence documents $d_{\neg c_{i}}$: (1) Both $d_{c_{i}}$ and $d_{\neg c_{i}}$ in the query language (); (2) $d_{c_{i}}$ in English, $d_{\neg c_{i}}$ in the query language (); (3) Both $d_{c_{i}}$ and $d_{\neg c_{i}}$ in English (); (4) $d_{c_{i}}$ in the query language, $d_{\neg c_{i}}$ in English (). Higher citation accuracy for variants and compared to and indicates that the model prefers citing documents in the query language. Conversely, higher accuracy for and suggests a persistent English preference regardless of the query language.
Results. We report citation accuracies for the four variants in Figure 4, broken down by target language for each model. Across more than half of the model-language combinations (28 out of 48), we observe the highest citation accuracy when the cited document is in the query language and all other documents are in English (). In 17 of these 28 cases, the second-best performance is when all documents are in the query language (). Since the configuration generally outperforms the variant, this suggest that models benefit from a language contrast between the cited and the remaining documents rather than simply having more documents match the query language. French follows this trend most strongly, with 4 out of 6 models exhibiting it. One possible explanation is that, as a relatively high-resource language, models have strong enough French representations, allowing them to effectively leverage the contrast and identify the most relevant document in context.
We further see that smaller models (8B) generally achieve lower accuracies than larger models (e.g., LLaMA-3.3 70B, Gemma-3 27B), extending our earlier observation from Section 5.1 that model size improves citation accuracy for English to non-English settings as well. Larger models also exhibit citation accuracies that are more tightly clustered across the four variants, suggesting greater robustness to language variation in the input context.
Together, these results suggest that query language plays a key role in models’ language preference: models tend to favor citing documents in the same language as the query, even when that language is not English. Interestingly, this mirrors findings in scientometrics literature, where humans also exhibit an “own-language preference”, tending to select and cite sources in the language of their writing (Yitzhaki, 1998; EGGHE et al., 2005). Detailed numerical results are in Appendix D.5.
7 Relevance vs. Language Preference
Sections 5 and 6 analyzed language preference in a controlled setup where all provided evidence documents were relevant to the query. In reality, however, retrievers are imperfect, and retrieved evidence often contains irrelevant or partially relevant documents (Chen et al., 2023; Jin et al., 2024). To better approximate such conditions, we relax the assumption that all documents are relevant and ask: between relevance and language, which exerts a stronger influence on model citation behavior?
Setting. We compare the effects of document relevance and language by varying the language of one relevant and one irrelevant document under three conditions: (1) En-En (                                                                                                                                                                                                                                         ): Both relevant and irrelevant documents are in English; (2) tgt-En (): Relevant document in the target language, irrelevant document in English; (3) En-tgt (): Relevant document in English, irrelevant document in the target language. Since ELI5 dataset does not include irrelevant documents, we use MIRACL (Zhang et al., 2023), a multilingual RAG dataset with Wikipedia queries. We use the English subset of the development set, restricting to queries with exactly one relevant document (231 queries). We randomly use one of the irrelevant documents. For each query, we follow the same process described in Section 3. Dataset statistics are in Appendix Table 2.
Figure 5: Accuracy per model with one relevant and one irrelevant evidence document in different languages. : Relevant doc in target language, irrelevant doc in English; : Relevant doc in English, irrelevant doc in target language;  : Baseline, both docs in English. Models trade off document relevance for language preference. Detailed results are in Appendix Figure 5.
Results. Our hypotheses are: (a) If citation accuracy in tgt-En is lower than the En-En baseline, it suggests the model is overly influenced by language, preferring to cite an irrelevant English document over a relevant target language one, and (b) If citation accuracy in En-tgt exceeds the baseline, it implies the model more easily ignores irrelevant target language distractors, again signaling English preference. As shown in Figure 5, results support both hypotheses. When the relevant document is in the target language, accuracies consistently drop below the baseline, indicating that irrelevant English content more easily mislead the model. Conversely, accuracies for all languages and models rise above the baseline for En-tgt, suggesting that target language distractors are easier to dismiss than English distractors. This aligns with recent findings that distractors in the same language as the relevant document degrade performance more severely (Qi et al., 2025). One interesting observation is Swahili. Despite yielding the lowest accuracies in ELI5 experiments (see Table 1), its performance in the En-tgt setup is relatively high. A possible explanation is its shared Latin script with English, which may make irrelevant Swahili documents appear more plausible choice. Full numerical results can be found in Appendix D.6.151515We further show that models also trade-off relevance for language preference when queries are posed in a language other than English in Appendix F.
8 Conclusion
We propose a controlled methodology to measure language preference in long-form mRAG by isolating language effects while controlling for document content and relevance. Our analysis shows that models preferentially cite English documents when queries are in English, with this bias stronger for lower-resource languages and mid-context. Importantly, this preference can outweigh relevance, with models often citing irrelevant English documents over relevant non-English ones. Overall, our findings demonstrate how model internals reveal citation behavior in mRAG and offer insights for designing more robust, inclusive systems that balance language and relevance.
Limitations. Our analysis uses a controlled setup with simplifying assumptions: (1) retrieval is complete and all evidence documents are equally relevant; (2) comparisons are pairwise with English; and (3) multilingual RAG is simulated via machine translations of English documents. [^16] We show our findings remain consistent with an alternative translation system in Appendix G. These assumptions may not hold in real-world settings, which could limit the generalizability of our results. Even so, our study provides valuable insights into language preference to guide future work on better understanding model citation behavior.
Ethics Statement
We recognize that research on language preference in multilingual retrieval-augmented generation can raise issues of fairness, particularly for low-resource and underrepresented languages. Our study aims to highlight such disparities rather than reinforce them. We caution that models may amplify English preference in multilingual contexts, with potential implications for inclusivity and equitable knowledge access.
Reproducibility Statement
All datasets used in our experiments are publicly available (ELI5, MIRACL) and are described in Section 4 and Appendix B, including dataset statistics, splits, and preprocessing steps. Our methodology for constructing multilingual evidence sets, generating English reference reports, and verifying citation correctness is documented in Section 3.1, with prompt templates provided in Appendix A. We report full per-language translation quality scores (Appendix D.1) and include statistical tests to assess the significance of results (Section 3.2, 5.1, and Appendix D). To support reproducibility, we will release our code and processed data upon publication, enabling others to replicate both our measurement pipeline and analyses.
Acknowledgements
We would like to thank the members of the Johns Hopkins University SCALE 2025 program. We are grateful to the generation team who gave constructive feedback and support in shaping this work. Dayeon also extends special thanks to the friends for making the internship experience in Baltimore truly memorable, including Yu Hou, Bryan Li, Gabrielle Kaili-May Liu, Maxime Dassen, Roxana Petcu, Jia-Huei Ju, Francois Landry, and Siddharth Singh. This work was supported in part by NSF Fairness in AI Grant 2147292, and by the Institute for Trustworthy AI in Law and Society (TRAILS), which is supported by the National Science Foundation under Award No. 2229885. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Appendix
Appendix A Prompts
We present the prompts used for generating the gold citation-supported report (Figure 6), obtaining supportedness judgments from LLM-as-judge (Figure 7), and guessing the next token predictions from the evaluated models (Figure 8). We adopt base prompts from GPTResearcher (Elovic, 2023).
Prompt A.1. Gold Report Generation Prompt Information: Document ID: {document ID} Title: {title} Content: {content} $\dots$ Using the above information, respond to the following query or task: {query}. The response should focus on the answer to the query, should be well structured, informative, and concise, with facts and numbers if available. Please follow all of the following guidelines in your response: - You MUST write in a single paragraph and at most {total words} words. - You MUST write the response in the following language: {language}. - You MUST cite your sources, especially for relevant sentences that answer the question. - When using information that comes from the documents, use citation which refer to the Document ID at the end of the sentence (e.g., [1]). - Do NOT cite multiple documents at the end of the sentence (e.g., [1][2]). - If multiple documents support the sentence, only cite the most relevant document. - It is important to ensure that the Document ID is a valid string from the information above and that the information in the sentence is present in the document. Response:
Figure 6: Prompt for generating gold citation-supported reports. Information section is populated with the document ID, title, and content of each evidence document. Boldface is only for emphasis.
Prompt A.2. LLM-as-judge Prompt Instruction: You are given a query, a document, and a sentence from a generated response that cites the document in answering the query. Determine which document best supports the information in the cited sentence. Respond only with the exact document ID. Do not provide any additional explanation. Query: {query} Information: Document ID: {document ID} Title: {title} Content: {content} $\dots$ Cited sentence: {statement} Response:
Figure 7: Prompt for getting supportedness judgments from LLM-as-judge. Information section is populated with the document ID, title, and content of each evidence document. Boldface is only for emphasis.
Prompt A.3. Next Token Prediction Prompt Information: Document ID: {document ID} Title: {title} Content: {content} $\dots$ Using the above information, the response is the answer to the query or task: {query} in a single sentence. You MUST cite the most relevant document by including only its Document ID in brackets at the end of the sentence (e.g., [Document ID]). Do NOT include any additional words inside or outside the brackets. Please output ONLY the number of the Document ID that is most relevant to the sentence. Response: {statement} [
Figure 8: Prompt for guessing the next token prediction. Information section is populated with the document ID, title, and content of each evidence document. Boldface is only for emphasis.
Appendix B Details of Dataset, Languages, and Models
We provide detailed statistics of the two long-form RAG datasets used in our experiments (ELI5 and MIRACL) in Table 2. The characteristics of the eight tested languages, including their language family, script, linguistic typology, and resource level, are summarized in Table 3. For the models, Table 4 includes their context window size, HuggingFace model identifier, and officially (un)supported languages. Lastly, Table 5 reports COMET-QE scores for each target language.
Dataset # Queries Avg. # Words ($q$) Avg. # Words ($t$) Avg. # Words ($d$) Avg. # Sent ($d$) Avg. # $d$ per $q$
ELI5 270 15.25 9.64 76.82 4.26 3.49
MIRACL 231 6.87 2.63 / 2.83 106.59 / 115.80 5.41 / 5.88 1.00 / 9.31
Table 2: Detailed statistics of long-form RAG datasets used. We report statistics for ELI5 (Explain Like I’m Five) and MIRACL. For MIRACL, statistics are shown as relevant / irrelevant documents. $q$: query; $t$: title, $d$: evidence document.
Language Family Language Script Synthesis Word Order Resource Level # Speakers # Wikipedia Size
Indo-European English Latin analytic SVO high 1,130M 5,758,285
French Latin fusional SVO high 398M 2,325,608
Spanish Latin fusional SVO high 592M 1,669,181
Russian Cyrillic fusional SVO mid 260M 1,476,045
Bengali Bengali fusional SOV low 337M 63,762
Sino-Tibetan Chinese Chinese analytic SVO high 1,350M 1,246,389
Koreanic Korean Hangul agglutinative SOV mid 128M 1,133,444
Afro-Asiatic Arabic Arabic fusional VSO mid 630M 656,982
Niger-Congo Swahili Latin agglutinative SVO low 83M 47,793
Table 3: Characteristics of tested languages. For each language, we show language family, script, linguistic typologies (synthesis and word order), and resource level measured by the number of speakers and Wikipedia articles (Zhang et al., 2023).
Model Context Window HuggingFace Model Identifier Supported Langs Unsupported Langs
LLaMA-3 8B 128K meta-llama/Llama-3.1-8B-Instruct en, es, fr ar, bn, ru, ko, sw, zh
LLaMA-3 70B 128K meta-llama/Llama-3.3-70B-Instruct en, es, fr ar, bn, ru, ko, sw, zh
Qwen-3 8B 33K Qwen/Qwen3-8B en, ar, bn, es, fr, ru, ko, sw, zh -
Qwen-3 14B 33K Qwen/Qwen3-14B en, ar, bn, es, fr, ru, ko, sw, zh -
Gemma-2 27B 128K google/gemma-3-27b-it en, ar, bn, es, fr, ru, ko, sw, zh -
Aya23 8B 8,192 CohereLabs/aya-23-8B en, ar, es, fr, ru, ko, zh bn, sw
Table 4: List of evaluated models. We report the context window size, HuggingFace model identifiers, and the officially supported languages during pretraining. Note: Supported language information is extracted from each model’s technical report. We use ISO 639-1 codes for languages. We use Qwen-3 series models with enable thinking=False mode.
Language COMET-QE($q$, $q^{\prime}$) COMET-QE($t$, $t^{\prime}$) COMET-QE($d$, $d^{\prime}$)
Arabic (ar) 0.752 0.541 0.511
Bengali (bn) 0.824 0.584 0.559
Spanish (es) 0.823 0.583 0.564
French (fr) 0.822 0.582 0.566
Korean (ko) 0.816 0.584 0.555
Russian (ru) 0.780 0.557 0.528
Swahili (sw) 0.769 0.544 0.516
Chinese (zh) 0.777 0.561 0.534
Table 5: COMET-QE scores by language. We evaluate the machine translation (MT) quality of non-English queries ($q$), titles ($t$), and evidence documents ($d$) in the ELI5 dataset. Apostrophe () indicates MT. Higher scores indicate better MT quality.
Appendix C Human Annotation
To validate the two-step automatic filtering process described in Section 3 for identifying supported statements, we conduct a small-scale human annotation study on 60 sampled statements. We stratify the sample into 30 “supported” statements (passing both the LLM-as-Judge and NLI entailment filters and included in the final statement pool) and 30 “unsupported” statements (failing one or both filters). We conducted a power analysis to justify our sample size. Using a $t$-test for 2 independent samples [^17] https://www.statsmodels.org/stable/generated/statsmodels.stats.power.TTestIndPower.html, we find that 26 statements per label group (supported and unsupported, total 52) are required to detect a minimum effect size of Cohen’s $d$ of 0.8 with a significance level of $\alpha$ of 0.05, and desired power of 0.8.
For each query $q$, statement $s_{i}$, and cited document $d_{c_{i}}$, we ask annotators: “How well is the statement supported by the provided document?” Responses are given on a five-point Likert scale from 5 (Definitely) to 1 (Not at all), using instructions similar to those provided when prompting the judge LLMs (Figure 7). Figure 9 shows the full instructions and an example provided to annotators.
We recruit six annotators from Prolific [^18] https://www.prolific.com/ who resides in the United States with first, primary, and fluent language as English. We compensate each with USD 8 (equivalent to USD 16/hour), totaling USD 56 including Prolific platform fees. Each annotator evaluates 30 statements (15 supported and 15 unsupported) presented in randomized order. Inter-annotator agreement is moderate, with a Krippendorff’s alpha of 0.559. The average rating for supported statements is 4.15 out of 5, while unsupported statements average 2.49 out of 5. These results indicate strong alignment between our automatic filtering process and human judgments of statement supportedness. Figure 10 plots the rating distribution for each label group.
(a) Task Instructions
(b) Example Annotation
Figure 9: Full instructions and example provided to human annotators. The annotation task was hosted on a custom-built website. Annotators first viewed a brief task instruction (a), then evaluate 30 statements, with an example shown in (b).
Figure 10: Rating distribution for each label group. We plot the distribution of 180 judgments collected during human annotation (90 supported and 90 unsupported statements). Results show that annotators can reliably distinguish supported from unsupported statements based on their ratings.
Appendix D Detailed Results
D.1 Machine Translation Quality
We evaluate machine translation (MT) quality for translated queries, titles, and evidence documents using COMET-QE scores. We do not perform any filtering based on these scores. Table 5 reports average scores by language, and Figure 11 shows full score distributions. We find little evidence that MT quality drives English preference. Document COMET-QE scores (last column of Table 5) are lowest for Arabic (0.511) and Swahili (0.516), while Bengali shows a relatively high score (0.559). Yet, citation accuracies (Table 1) show that Arabic’s ranking varies widely across models – third lowest for LLaMA-3.1 8B and Qwen-3 8B, lowest for LLaMA-3.3 70B, fourth lowest for Qwen-3 14B and Aya23, but relatively higher for Gemma-3 27B. By contrast, Bengali exhibits the second-strongest English preference after Swahili despite its higher MT quality. These results suggest that resource level, rather than MT quality, is a stronger indicator of English preference.
(a) Arabic (ar)
(b) Bengali (bn)
(c) Spanish (es)
(d) French (fr)
(e) Korean (ko)
(f) Russian (ru)
(g) Swahili (sw)
(h) Chinese (zh)
Figure 11: COMET-QE score distributions by language. Distributions are more skewed for shorter content (e.g., title), while broader distributions for longer content (e.g., evidence document).
D.2 Evidence of English Preference
Language LLaMA-3.1 8B LLaMA-3.3 70B Qwen-3 8B Qwen-3 14B Gemma-3 27B Aya23 8B
English 0.651 0.991 0.758 0.984 0.980 0.527
Arabic 0.629 (-0.022) 0.990 (-0.001) 0.751 (-0.007) 0.979 (-0.005) 0.968 (-0.012) 0.463 (-0.064)
Bengali 0.647 (-0.004) 0.990 (-0.001) 0.736 (-0.022) 0.981 (-0.003) 0.977 (-0.003) 0.442 (-0.085)
Spanish 0.626 (-0.025) 0.987 (-0.004) 0.752 (-0.006) 0.981 (-0.003) 0.979 (-0.001) 0.483 (-0.044)
French 0.649 (-0.002) 0.991 (0.000) 0.728 (-0.030) 0.983 (-0.001) 0.973 (-0.007) 0.499 (-0.028)
Korean 0.620 (-0.031) 0.982 (-0.009) 0.730 (-0.028) 0.983 (-0.001) 0.955 (-0.025) 0.494 (-0.033)
Russian 0.634 (-0.017) 0.990 (-0.001) 0.707 (-0.051) 0.982 (-0.002) 0.961 (-0.019) 0.465 (-0.062)
Swahili 0.630 (-0.021) 0.987 (-0.004) 0.634 (-0.124) 0.967 (-0.017) 0.966 (-0.014) 0.479 (-0.048)
Chinese 0.642 (-0.009) 0.988 (-0.003) 0.706 (-0.052) 0.984 (0.000) 0.976 (-0.004) 0.488 (-0.039)
Table 6: Next token probabilities for the correct citation ID by model and language (↑). We present mean values along with the difference from English baseline indicated in subscript.
Language LLaMA-3.1 8B LLaMA-3.3 70B Qwen-3 8B Qwen-3 14B Gemma-3 27B Aya23 8B
English 1.106 0.132 0.388 0.064 0.028 1.215
Arabic 1.146 (+0.040) 0.176 (+0.044) 0.500 (+0.112) 0.088 (+0.024) 0.063 (+0.035) 1.277 (+0.062)
Bengali 1.169 (+0.063) 0.178 (+0.046) 0.457 (+0.069) 0.095 (+0.031) 0.051 (+0.023) 1.350 (+0.135)
Spanish 1.152 (+0.046) 0.150 (+0.018) 0.460 (+0.072) 0.081 (+0.017) 0.048 (+0.020) 1.260 (+0.045)
French 1.122 (+0.016) 0.149 (+0.017) 0.389 (+0.001) 0.075 (+0.011) 0.051 (+0.023) 1.247 (+0.032)
Korean 1.150 (+0.044) 0.166 (+0.034) 0.394 (+0.006) 0.087 (+0.023) 0.059 (+0.031) 1.269 (+0.054)
Russian 1.134 (+0.028) 0.162 (+0.030) 0.412 (+0.024) 0.074 (+0.010) 0.059 (+0.031) 1.266 (+0.051)
Swahili 1.194 (+0.088) 0.182 (+0.050) 0.508 (+0.120) 0.123 (+0.059) 0.054 (+0.026) 1.254 (+0.039)
Chinese 1.130 (+0.024) 0.159 (+0.027) 0.385 (+-0.003) 0.084 (+0.020) 0.067 (+0.039) 1.255 (+0.040)
Table 7: Shannon entropy by model and language (↓). We present mean values along with the difference from English baseline indicated in subscript.
In Tables 6 and 7, we report, for each model and language, (a) the next token probability assigned to the correct citation ID and and (b) the Shannon entropy of the next token distribution. Across all models, we observe consistently higher probabilities when the cited evidence document is in English, alongside lower entropy values. Together, this suggests that models are not only more accurate but also more confident when correctly citing English documents.
D.3 Position-wise Accuracy per Language
We show accuracy gap between English and each target language in Figure 12. We show that the findings with the aggregated results in Section 5.1 are consistent for all languages: the accuracy drop is generally most pronounced when the cited document appears in the middle of the input context.
(a) Arabic (ar)
(b) Bengali (bn)
(c) Spanish (es)
(d) French (fr)
(e) Korean (ko)
(f) Russian (ru)
(g) Swahili (sw)
(h) Chinese (zh)
Figure 12: Accuracy difference between English and each target language binned by relative position. Each bin is normalized by sample size.
D.4 Logit Lens Analysis
Figures 13 to 17 present logit lens visualizations for each model. We observe different trends:
LLaMA-3.3 70B. The model follows a trajectory similar to LLaMA-3.1 8B. Both the correct and wrong citation ID predictions begin to rise around layer 40, peak sharply at layers 52-57, then decline until layer 60 before increasing again and stabilizing toward the final layers. Throughout, correct predictions consistently outnumber incorrect ones. As with LLaMA-3.1 8B, the gap between correct and incorrect predictions narrows for lower-resource languages.
Qwen-3 8B. The model exhibits a staggered pattern, where correct citation IDs peak around layer 26, again at layers 28-30, and once more at the final layer, remaining low in between. While the model already predicts the correct IDs in earlier layers (28-30), they are overtaken by invalid predictions just before the final two layers, after which the model uncovers and ends with a final peak in accuracy.
Qwen-3 14B. Despite belonging to the same Qwen-3 family, this model exhibits a completely different behavior from Qwen-3 8B. For most of its layers, it fails to predict outputs in the expected citation format. Only in the final layers (38-40), we observe an increase in correct citation predictions, consistently outpacing incorrect ones. This suggests a more conservative prediction strategy, where it delays citation prediction until the very end, or it can only recognize the citation format at the final layers.
Gemma-3 27B. Similar to the Qwen-3 8B, this model shows a staggered pattern, where incorrect predictions remain low, while correct predictions generally increase. There are sharp drops around layers 53-54 and layer 58. However, the model recovers by the final layer, and the count of correct predictions stays high.
Aya23 8B. This model stands out from the others, as incorrect predictions generally outnumber correct ones. This aligns with the results in Table 1, where Aya23 8B shows the largest average accuracy drop for target languages. It is also especially pronounced for lower-resource languages like Bengali or Swahili, where the gap between correct and incorrect predictions is even wider.
Figure 13: Logit lens visualization per language for LLaMA-3.3 70B (80 layers). $x$-axis: Last layer index; $y$-axis: Statement count. We show the last 40 layers. : Correct citation ID of document in target language; : Wrong citation ID of document in English; : Not in valid citation set.
Figure 14: Logit lens visualization per language for Qwen-3 8B (36 layers). $x$-axis: Last layer index; $y$-axis: Statement count. We show the last 14 layers.
Figure 15: Logit lens visualization per language for Qwen-3 14B (40 layers). $x$-axis: Last layer index; $y$-axis: Statement count. We show the last 14 layers.
Figure 16: Logit lens visualization per language for Gemma-3 27B (62 layers). $x$-axis: Last layer index; $y$-axis: Statement count. We show the last 18 layers to capture the entire pattern.
Figure 17: Logit lens visualization per language for Aya23 8B (32 layers). $x$-axis: Last layer index; $y$-axis: Statement count. We show the last 14 layers.
D.5 Query Language Variants
In Table 8, we report the full numerical results when the query is posed in a target language. We consider four variants, differing in the language of the cited document and the remaining evidence documents, following the same notation introduced in Figure 4: (1) All tgt: all documents in the query language, (2) All tgt + Correct En: cited document in English and all other documents in the query language, (3) All En: all documents in English, and (4) All En + Correct tgt: cited document in the query language and all other documents in English. Overall, we find that models tend to prefer citing evidence in the query language, with All En + Correct tgt configuration achieving the highest accuracy in more than half of the cases.
Model Language All tgt () All tgt + Correct En () All En () All En + Correct tgt ()
LLaMA-3.1 8B Arabic 0.616 0.572 0.557 0.658
Bengali 0.673 0.619 0.683 0.729
Spanish 0.683 0.639 0.667 0.717
French 0.716 0.713 0.674 0.716
Korean 0.645 0.646 0.611 0.667
Russian 0.736 0.666 0.686 0.745
Swahili 0.627 0.638 0.640 0.648
Chinese 0.607 0.631 0.579 0.610
LLaMA-3.3 70B Arabic 0.828 0.843 0.858 0.837
Bengali 0.883 0.878 0.875 0.890
Spanish 0.875 0.886 0.877 0.883
French 0.875 0.893 0.880 0.906
Korean 0.866 0.893 0.880 0.857
Russian 0.893 0.912 0.900 0.901
Swahili 0.902 0.902 0.904 0.879
Chinese 0.792 0.815 0.772 0.811
Qwen-3 8B Arabic 0.635 0.598 0.583 0.626
Bengali 0.605 0.560 0.650 0.632
Spanish 0.648 0.600 0.603 0.665
French 0.621 0.575 0.585 0.650
Korean 0.677 0.705 0.672 0.655
Russian 0.710 0.686 0.648 0.673
Swahili 0.477 0.459 0.463 0.479
Chinese 0.538 0.533 0.554 0.487
Qwen-3 14B Arabic 0.832 0.773 0.787 0.843
Bengali 0.910 0.892 0.901 0.909
Spanish 0.877 0.842 0.845 0.906
French 0.883 0.857 0.868 0.908
Korean 0.858 0.843 0.853 0.843
Russian 0.889 0.867 0.875 0.898
Swahili 0.759 0.735 0.743 0.697
Chinese 0.744 0.765 0.737 0.730
Gemma-3 27B Arabic 0.823 0.808 0.814 0.859
Bengali 0.868 0.867 0.873 0.897
Spanish 0.852 0.854 0.852 0.897
French 0.863 0.845 0.861 0.898
Korean 0.844 0.862 0.853 0.867
Russian 0.878 0.851 0.867 0.902
Swahili 0.832 0.806 0.836 0.896
Chinese 0.811 0.792 0.779 0.792
Aya23 8B Arabic 0.464 0.443 0.475 0.516
Bengali 0.454 0.401 0.354 0.537
Spanish 0.563 0.577 0.555 0.569
French 0.551 0.574 0.575 0.578
Korean 0.574 0.537 0.501 0.572
Russian 0.531 0.540 0.532 0.590
Swahili 0.427 0.312 0.358 0.528
Chinese 0.453 0.460 0.492 0.488
Table 8: Numerical results when the query is in target language. We report accuracies for four variants per model and language. We use the same shape notation as in Figure 4. Best scores for each row is bold.
D.6 Relevance vs. Language Preference
In Figure 18, we plot citation accuracy for the remaining models (Qwen-3 8B and Aya23 8B) with one relevant and one irrelevant evidence document in different languages, complementing the results in Figure 5. Table 9 reports the full numerical results using the notation from Section 7: (1) En-En: both relevant and irrelevant documents are in English, (2) tgt-En: relevant document in the target language and irrelevant document in English, and (3) En-tgt: relevant document in English and irrelevant document in the target language. Overall, we observe that citation accuracy in tgt-En is generally lower than the En-En baseline, while En-tgt is consistently higher, both indicating a strong English preference that persists regardless of differences in document relevance.
Figure 18: Accuracy per model with one relevant and one irrelevant evidence document in different languages. : Relevant doc in target language, irrelevant doc in English; : Relevant doc in English, irrelevant doc in target language;  : Baseline, both docs in English.
Model Language En-En tgt-En (↓) En-tgt (↑)
LLaMA-3.1 8B Arabic 0.944 0.931 0.961
Bengali 0.944 0.918 0.965
Spanish 0.944 0.913 0.970
French 0.944 0.939 0.970
Korean 0.944 0.952 0.974
Russian 0.944 0.965 0.961
Swahili 0.944 0.974 0.961
Chinese 0.944 0.944 0.970
LLaMA-3.3 70B Arabic 0.974 0.935 0.983
Bengali 0.974 0.944 0.987
Spanish 0.974 0.926 0.987
French 0.974 0.957 0.987
Korean 0.974 0.944 0.978
Russian 0.974 0.957 0.987
Swahili 0.974 0.961 0.978
Chinese 0.974 0.952 0.983
Qwen-3 8B Arabic 0.831 0.796 0.948
Bengali 0.831 0.766 0.931
Spanish 0.831 0.714 0.870
French 0.831 0.775 0.883
Korean 0.831 0.836 0.961
Russian 0.831 0.805 0.909
Swahili 0.831 0.710 0.913
Chinese 0.831 0.771 0.922
Qwen-3 14B Arabic 0.961 0.926 0.970
Bengali 0.961 0.918 0.987
Spanish 0.961 0.922 0.974
French 0.961 0.944 0.970
Korean 0.961 0.918 0.974
Russian 0.961 0.935 0.970
Swahili 0.961 0.896 0.974
Chinese 0.961 0.918 0.961
Gemma-3 27B Arabic 0.944 0.887 0.970
Bengali 0.944 0.905 0.970
Spanish 0.944 0.862 0.974
French 0.944 0.905 0.961
Korean 0.944 0.905 0.965
Russian 0.944 0.931 0.952
Swahili 0.944 0.887 0.961
Chinese 0.944 0.883 0.965
Aya23 8B Arabic 0.736 0.753 0.840
Bengali 0.736 0.623 0.844
Spanish 0.736 0.719 0.818
French 0.736 0.740 0.827
Korean 0.736 0.714 0.827
Russian 0.736 0.714 0.827
Swahili 0.736 0.567 0.866
Chinese 0.736 0.727 0.792
Table 9: Numerical results for setup with one relevant and one irrelevant evidence document, in different languages. We use the same label as in Figure 18. Red denotes tgt-En scores that are lower than the En-En baseline; Green denotes En-tgt scores that are higher than the baseline.
Appendix E Contributive Attribution Patterns
Our analysis of language preference has been based on corroborative attribution, measuring the probability of generating in-line citations, which identifies sources that support a statement (Menick et al., 2022; Liu et al., 2023). However, if models are citing more English documents, that does not necessarily mean they are actually attributing on their content. If models truly favor English sources, we would expect that preference to also appear when we examine contributive attribution, which identifies sources that cause a model to generate a specific statement.
Figure 19: Hit@1 and Score@1 by model and language. Higher values indicate more accurate attribution to the cited document.
To test this, we use an attribution model ContextCite (Cohen-Wang et al., 2024), which estimates the influence of each document on the model’s generation. ContextCite is a fitted linear surrogate model that encodes the importance of each source in the context by taking ablated contexts $m\in\{0,1\}^{K}$ as input, where $m_{j}=1$ indicates that sentence $j$ is present and $m_{j}=0$ indicates that it is masked. The model predicts the ground-truth logit-scaled probability for a given mask $m$ as:
$f(m)=w^{\intercal}m+b,$ (4)
where $w\in\mathbb{R}^{K}$ contains per-sentence attribution weights and $b$ is a bias term.
In our case, given a query $q$, a set of $K$ relevant documents $\mathcal{D}=\{d_{1},\dots,d_{k}\}$, and a pool of statements $\{s_{i}\}$, ContextCite returns a ranked list of sentences from $\mathcal{D}$ that most influenced the generation of each $s_{i}$, along with their attribution scores. Here, $\mathcal{D}$ is composed of the cited document in the target language and all remaining documents in English. We evaluate attribution quality using two metrics: (1) Hit@1 (↑): whether the top-ranked sentence originates from the cited document, and (2) Score@1 (↑): the attribution score $w_{j^{*}}$ of the top-ranked sentence, indicating its estimated relative importance to the model’s prediction.
Figure 19 presents both metrics by each model and language. Across all models, both metrics peak when the cited document is in English, outperforming all target language counterparts. This suggests that English preference is not merely a surface-level citation pattern but reflects more reliance on English sources during generation. Full numerical results for Hit@$k$ and Score@$k$ ($k\in\{1,3\}$) are presented in Table 12.
Appendix F Relevance vs. Query Language Preference
We conduct the same set of experiments from Section 7 in a setting where the query is in a language other than English. We use the same dataset, MIRACL (231 queries) with machine translated queries, relevant, and irrelevant documents. While fixing the query in target language, we vary the language of one relevant and one irrelevant document under three conditions: (1) tgt-tgt (): Both relevant and irrelevant documents are in the target (query) language; (2) tgt-En (): Relevant document in the target language, irrelevant document in English; (3) En-tgt (): Relevant document in English, irrelevant document in the target language.
Our hypotheses are: (a) If citation accuracy in En-tgt is lower than tgt-tgt or tgt-En, it suggests that models trade off relevance for query language preference, citing irrelevant target language documents over relevant English ones, and (b) If citation accuracy of tgt-En exceeds tgt-tgt, it indicates that models more easily ignore irrelevant English distractors, showing stronger query language preference over English preference.
As shown in Figure 20, results support the first hypothesis: citation accuracies are generally the lowest when the relevant document is in English and the distractor is in the query language, showing that models preferring language over relevance persists for non-English queries. Interestingly, we show that this trend is most evident for (i) lower-resource languages such as Bengali (bn) and Swahili (sw) and (ii) models that reported to support all tested target languages (Qwen-3 8B and 14B, Gemma-3 27B; Table 4). Conversely, we show mixed results for the second hypothesis. The citation accuracies of tgt-En and tgt-tgt are largely similar, suggesting that English distractors are not necessarily easier at misleading models than those in the target language. Overall, our results imply that models show a consistent preference for the query language over relevance, but the distractor’s language matters less when the query is not in English.
Figure 20: Accuracy per model with one relevant and one irrelevant evidence document in different languages. : Both docs in target language; : Relevant doc in target language, irrelevant doc in English; : Relevant doc in English, irrelevant doc in target language. Models also trade off document relevance for language preference for queries not in English.
Appendix G Alternative Machine Translation System
We replicate the main experiments from Section 5 using an alternative machine translation (MT) system to verify whether the English preference trend persists. Specifically, we use Tower-Instruct 7B [^19] Unbabel/TowerInstruct-7B-v0.2 (Alves et al., 2024), a model trained for diverse translation-related tasks including general MT, automatic post-editing, and grammatical error correction. Table 10 reports COMET-QE scores by language. Compared to Google Translate (see Table 5), MT quality shows mixed results, with higher scores for all languages except Arabic and Bengali.
Table 11 presents citation accuracies per model and language using Tower-Instruct translations. We show that the general trend observed with Google Translate persists: models achieve the highest citation accuracy when the cited document is in English. The accuracy gaps between English and target languages are all statistically significant ($p<$ 0.001), despite using a stronger MT system compared to Google Translate. This suggests that the English preference cannot be fully attributed to using machine-translated documents. Notably, citing Arabic documents leads to the largest performance drop relative to English across all models, likely reflecting the lower COMET-QE scores for Arabic shown in Table 10.
Appendix H Usage of Large Language Models
We used LLMs to support and refine the writing of our work. Importantly, we did not rely on them to generate content or sentences from scratch. Instead, we employed them primarily to polish the clarity and expression of how we presented our results. In addition, we used them for stylistic adjustments, such as improving readability and removing layout issues (e.g., widows and orphans).
Language COMET-QE($q$, $q^{\prime}$) COMET-QE($t$, $t^{\prime}$) COMET-QE($d$, $d^{\prime}$)
Arabic (ar) 0.467 0.374 0.311
Bengali (bn) 0.802 0.562 0.491
Spanish (es) 0.855 0.595 0.574
French (fr) 0.859 0.598 0.548
Korean (ko) 0.857 0.597 0.554
Russian (ru) 0.839 0.588 0.549
Swahili (sw) 0.787 0.549 0.482
Chinese (zh) 0.817 0.574 0.505
Table 10: COMET-QE scores by language using Tower-Instruct 7B translations. We evaluate the machine translation (MT) quality of non-English queries ($q$), titles ($t$), and evidence documents ($d$) in the ELI5 dataset. Apostrophe () indicates MT. Higher scores indicate better MT quality.
Language LLaMA-3.1 8B LLaMA-3.3 70B Qwen-3 8B Qwen-3 14B Gemma-3 27B Aya23 8B
English 67.4 85.9 62.6 83.0 86.2 60.0
Arabic 24.6 (-42.8) 21.1 (-64.8) 15.9 (-46.7) 26.3 (-56.7) 26.4 (-59.8) 23.2 (-36.8)
Bengali 45.5 (-22.0) 52.9 (-33.0) 34.1 (-28.5) 53.4 (-29.6) 52.4 (-33.8) 38.6 (-21.4)
Spanish 58.5 (-8.91) 72.5 (-13.4) 50.2 (-12.4) 73.4 (-9.64) 74.2 (-12.0) 47.5 (-12.5)
French 53.0 (-14.4) 65.7 (-20.2) 41.8 (-20.8) 66.4 (-16.6) 65.0 (-21.2) 42.5 (-17.4)
Korean 55.2 (-12.2) 62.6 (-23.3) 45.5 (-17.1) 65.7 (-17.3) 69.8 (-16.4) 41.0 (-19.0)
Russian 55.1 (-12.3) 71.6 (-14.3) 48.9 (-13.7) 68.1 (-14.9) 70.1 (-16.1) 43.3 (-16.7)
Swahili 41.7 (-25.7) 49.5 (-36.4) 34.2 (-28.4) 50.9 (-32.1) 48.9 (-37.3) 37.1 (-22.9)
Chinese 44.2 (-23.2) 55.8 (-30.1) 37.4 (-25.2) 57.7 (-25.3) 54.8 (-31.4) 39.3 (-20.7)
Table 11: Citation accuracies (%) by model and language using Tower-Instruct 7B translations. We present mean accuracy values $\mathrm{\textbf{Acc}}^{(\ell)}$ along with $\Delta(\ell_{\mathrm{target}})$ in subscript as percent (%). Pairwise two-sided $t$-tests are performed to compare accuracy between English and the target language, with the null hypothesis that the mean citation accuracy is equal across languages. Bonferroni correction is applied for multiple comparisons. All differences are statistically significant ($p$ << 0.001). Color coding indicates the magnitude of $\Delta(\ell_{\mathrm{target}})$: largest, second largest, others.
Model Language Hit@1 (↑) Hit@3 (↑) Score@1 (↑) Score@3 (↑)
LLaMA-3.1 8B English 0.880 0.971 10.737 11.114
Arabic 0.771 0.928 7.370 7.777
Bengali 0.777 0.908 8.648 9.041
Spanish 0.821 0.934 8.797 9.195
French 0.824 0.943 8.641 9.051
Korean 0.758 0.924 6.649 7.064
Russian 0.804 0.953 8.045 8.495
Swahili 0.718 0.903 5.768 6.197
Chinese 0.821 0.929 9.976 10.418
LLaMA-3.3 70B English 0.910 0.968 14.749 13.080
Arabic 0.837 0.955 10.138 10.594
Bengali 0.837 0.943 12.874 13.369
Spanish 0.851 0.970 11.081 11.555
French 0.860 0.970 10.918 11.401
Korean 0.832 0.960 9.249 9.695
Russian 0.861 0.971 10.647 11.117
Swahili 0.773 0.937 8.363 8.927
Chinese 0.875 0.984 12.576 15.075
Qwen-3 8B English 0.881 0.966 13.128 13.563
Arabic 0.693 0.866 7.183 7.634
Bengali 0.674 0.826 7.912 8.376
Spanish 0.769 0.932 9.200 9.762
French 0.779 0.910 9.274 9.768
Korean 0.684 0.865 6.642 7.138
Russian 0.768 0.929 8.602 9.154
Swahili 0.444 0.711 3.042 3.487
Chinese 0.755 0.862 10.703 11.053
Qwen-3 14B English 0.880 0.973 12.986 13.399
Arabic 0.746 0.894 8.054 8.472
Bengali 0.722 0.881 9.197 9.683
Spanish 0.818 0.957 10.075 10.562
French 0.826 0.949 10.031 10.465
Korean 0.740 0.899 7.393 7.877
Russian 0.801 0.941 9.227 9.685
Swahili 0.495 0.745 3.918 4.396
Chinese 0.785 0.894 11.511 12.073
Gemma-3 27B English 0.902 0.976 12.321 12.752
Arabic 0.704 0.910 6.891 7.253
Bengali 0.535 0.722 5.784 5.781
Spanish 0.781 0.938 8.347 8.853
French 0.806 0.932 8.575 9.029
Korean 0.727 0.903 6.516 7.128
Russian 0.776 0.911 8.223 8.771
Swahili 0.602 0.657 8.002 2.256
Chinese 0.751 0.898 9.504 9.955
Aya23 8B English 0.862 0.963 11.729 12.247
Arabic 0.737 0.912 7.180 7.671
Bengali 0.423 0.683 2.087 2.600
Spanish 0.777 0.932 8.665 9.210
French 0.804 0.927 8.811 9.280
Korean 0.729 0.919 6.622 7.143
Russian 0.765 0.918 8.049 8.608
Swahili 0.326 0.634 1.627 1.990
Chinese 0.760 0.883 9.955 10.443
Table 12: Numerical results for ContextCite. We report Hit@$k$ and Score@$k$ (where $k\in\{1,3\}$) for each model across all languages.